OpenStudy (anonymous):

please help!!!! for medal!! The function f(x) = 16(2)x represents the growth of a bee population every year in a remote swamp. Jennifer wants to manipulate the formula to an equivalent form that calculates two times a year, not just once a year. Which function is correct for Jennifer's purpose, and what is the new growth rate?

OpenStudy (anonymous):

@phi i have no idea what to do

OpenStudy (phi):

maybe we write it as \[ f(x)= 16\cdot 2^x = 16 \cdot (1+1)^x \] and to change to every 6 months \[ f(x)= 16\cdot (1+1/2)^{2x}= 16 \cdot 1.5^{2x} \] but I am guessing

OpenStudy (anonymous):

f(x) = 16(2)x; growth rate 200% f(x) = 16(2)2x; growth rate 8% f(x) = 16(1.41)x; growth rate 8% f(x) = 16(1.41)2x; growth rate 41% these are the choices

OpenStudy (anonymous):

if this helps you @phi

OpenStudy (phi):

I would go with choice D

OpenStudy (anonymous):

okayy agian thank you so much !

OpenStudy (phi):

I guess the idea is to write the exponent as 2x (twice as many as before) but to not change things we do this \[ \left(2\right)^{\frac{1}{2} \cdot 2x } \] and then interpret that as \[ \left(2^{\frac{1}{2} }\right)^{2x} \] and 2^(1/2) is sqr(2) = 1.41, so \[ (1.41)^{2x} \] and we put the 16 in there for the final answer \[ 16(1.41)^{2x} \]

OpenStudy (anonymous):

thank you!